Abby purchased 100 shares of her father’s favorite stock for $27 per share exactly 1 year ago, commission free. She sold it today for a total amount of $3200. She plans to invest the entire amount in a different corporation’s stock today, but she must now pay a $30 commission fee. If she plans to sell this new stock exactly 1 year from now and realize the same return as she has just made, what must be the total amount she receives next year? Include the commission fee as a part of the purchase price, but neglect any tax effects. The total amount that she receives next year is $____.
Answer #1

Total amount Abby invested exactly year ago = 27$*100 = 2700

Total amount Abby got after selling the shares = 3200

Rate of return = ((Ending value of investment/ Beginning value of investment)-1)*100

ROR = ((3200/2700)-1)*100

ROR = (1.185-1)*100

ROR = 18.5%

Now new stock purchase price = Commission fee + Total amount after selling the old shares

New stock purchase price = 30+3200 = 3230

ROR = 18.5%

Amount he should get after one year = X

ROR = ((Ending value/Beginning value)-1)*100

18.5 = ((X/3230)-1)*100

0.185 = (X/3230)-1

3230 (1+0.185) = X

X = 3827.55$

So the total amount she receives next year = 3827.55$